The electric current in a charging R-C circuit is given by i = i0 e-t/RC where i0, R and C are constant parameters of current at (a) t = 0, (b) t = RC, (c) t = 10 RC.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. Given that, i = i0e-t/RC ∴ Rate of change of current = di/dt = d/dt i0e-i/RC = i0 d/dt e-t\RC = -i_0/RC x e-t/RC When (a) t = 0, di/dt = (-i)/RC (b) when t = RC, di/dt = (-i)/RCe (c) when t = 10 RC, di/dt = 〖-i〗_0/〖RCe〗^10